Find the linear transformation to reducethe following quadratic forms into cononical form using diagonalation method 3x^2+3y^2+3z^2+2xy+2xz-2yz
1
Expert's answer
2021-09-09T14:27:48-0400

Solution;

The matrix representation,A,of the equation is;

"A=\begin{pmatrix}
 3 & 1&1\\
 1 & 3&-1\\
1&-1&3


\end{pmatrix}"

Find the characteristic equation from,

"|A-\lambda I|=0"

i.e;

"\lambda^3-s_1\lambda^2+s_2\lambda-s_3=0"

In which ;

"s_1" =Sum of the main diagonal elements.

"s_1=3+3+3=9"

"s_2" = Sum of the minors of the main diagonal elements;

"s_2=\begin{vmatrix}
 3 & 1 \\
 1& 3
\end{vmatrix}" +"\begin{vmatrix}
 3 & 1\\
 1 & 3
\end{vmatrix}" +"\begin{vmatrix}
 3& -1 \\
 -1 & 3
\end{vmatrix}"

"s_2=24"

"s_3" =|A|="\begin{vmatrix}
 3& 1&1 \\
 1 & 3&-1\\
1&-1&3
\end{vmatrix}" =16

The characteristic equation is ;

"\lambda^3-9\lambda^2+24\lambda-16=0"

Solve for the roots of the equation;

If "\lambda=1"

1-9+24-16=0

Hence ,"\lambda =1" is a root.

By synthetic Division,other roots are from;

"\lambda ^2-8\lambda+16=0"

"(\lambda-4)(\lambda-4)=0"

"\lambda=4,4"

Hence the eigenvalues are;

4,4,1

Find the eigenvectors;

(A-"\lambda I" )x=0

"[\begin{pmatrix}
 3 &1&1\\
 1& 3&-1\\
1&-1&3
\end{pmatrix}" -"\begin{pmatrix}
 \lambda & 0 \\
 0 & \lambda&0\\
0&0&\lambda
\end{pmatrix}]x=0"

To have;

"\begin{bmatrix}
 3-\lambda & 1&1 \\
 1 & 3-\lambda&-1\\
1&-1&3-\lambda
\end{bmatrix}" "\begin{bmatrix}
 x \\
 y\\
z\\
\end{bmatrix}=0"

Take;

Case (i),if "\lambda" =4,the equation becomes;

"\begin{bmatrix}
 -1& 1&1\\
 1 & -1&-1\\
1&-1&-1
\end{bmatrix}" "\begin{bmatrix}
 x \\
 y\\
z\\
\end{bmatrix}=0"

-x+y+z=0

x-y-z=0

x-y-z=0

The equation gives ;

"x_1=\begin{bmatrix}
 2\\
 1\\
1
\end{bmatrix}"

Case(iii),"\lambda =1" ,the equation becomes;

"\begin{bmatrix}
 2 & 1&1\\
 1 & 2&-1\\
1&-1&2
\end{bmatrix}" "\begin{bmatrix}
 x\\
 y\\
z
\end{bmatrix}=0"

2x+y+z=0

x+2y-z=0

x-y+2y=0

The equations give;

"x_3=\begin{bmatrix}
 -1\\
 1\\
1
\end{bmatrix}"

To find the third set of eigenvectors,"x_3"

Since;

"x_1x_2^T=0"

"x_2x_3^T=0"

"x_3x_1^T=0"

We know;

"x_1=\begin{bmatrix}
 2 \\
 1\\
1
\end{bmatrix}" ,"x_3=\begin{bmatrix}
 -1\\
 1\\
1
\end{bmatrix}"

Let ;

"x_2=\begin{bmatrix}
 a \\
 b\\
c
\end{bmatrix}"

Hence;

"\begin{bmatrix}
 2 & 1&1 \\
 -1 & 1&1
\end{bmatrix}""\begin{bmatrix}
 a \\
 b\\
c
\end{bmatrix}=0"

2a+b+c=0

-a+b+c=0

Hence,

"x_2=\begin{bmatrix}
 0\\
 -1\\
1\\
\end{bmatrix}"

The model matrix,M from the eigenvectors is;

"\begin{pmatrix}
 2 & 0&-1\\
 1 & -1&1\\
1&1&1
\end{pmatrix}"

The normalised model matrix ,N,is as follows;

Length of eigenvectors;

"x_1=\sqrt{4+1+1}=\sqrt6"

"x_2=\sqrt{0+1+1}=\sqrt2"

"x_3=\sqrt{1+1+1}=\sqrt 3"

"N=\begin{bmatrix}
 \frac{2}{\sqrt6}&0&\frac{-1}{\sqrt3} \\
 \frac{1}{\sqrt6} & \frac{-1}{\sqrt2}&\frac{1}{\sqrt3}\\
\frac{1}{\sqrt6}&\frac{1}{\sqrt2}&\frac{1}{\sqrt3}
\end{bmatrix}"

Transponse normalised model matrix ,"N^T" ;

"N^T=\begin{bmatrix}
 \frac{2}{\sqrt6}&\frac{1}{\sqrt6}&\frac{1}{\sqrt6} \\
 0 & \frac{-1}{\sqrt2}&\frac{1}{\sqrt2}\\
\frac{-1}{\sqrt3}&\frac{1}{\sqrt3}&\frac{1}{\sqrt3}
\end{bmatrix}"

Find the diagonalized matrix,"D=N^TAN"

"D=\begin{bmatrix}
 4 & 0&0 \\
 0 & 4&0\\
0&0&1
\end{bmatrix}"

Hence D is the matrix of linear transformation of the quadratic form into canonical form.

Thus the linear transformation,X is;

X=Dy

X="\begin{pmatrix}
 4& 0&0 \\
 0 & 4&0\\
0&0&1
\end{pmatrix}" "(y_1,y_2,y_3)"